# Optimal Modified Feedback Strategies in LQ Games under Control Imperfections

### v2

@@ Abstract
Game-theoretic approaches and Nash equilibrium have been widely applied across various engineering domains. However, practical challenges such as disturbances, delays, and actuator limitations can hinder the precise execution of Nash equilibrium strategies. This work investigates the impact of such implementation imperfections on game trajectories and players’ costs in the context of a two-player finite-horizon linear–quadratic (LQ) nonzero-sum game. Specifically, we analyze how small deviations by one player, measured or estimated at each stage, affect the state and cost function of the other player. To mitigate these effects, we propose an adjusted control policy that optimally compensates for the deviations under the stated information structure and can, under certain conditions, exploit them to improve performance. Rigorous mathematical analysis and proofs are provided, and the effectiveness of the proposed method is demonstrated through a representative numerical example.
@@ I Introduction
Game theory is a powerful analytical framework for modeling rational decision-making in complex multi-agent systems, particularly when interactions are strategic and agents’ decisions influence each other [1, 2]. Such interactions often involve conflicting objectives, motivating the development of solution concepts to characterize equilibrium behavior. Among these, the Nash equilibrium is a well-established notion, defined as a set of strategies from which no participant can unilaterally deviate to achieve a better outcome [3].
Thanks to its solid theoretical foundation, Nash equilibrium has been widely applied in diverse engineering domains, including robotics [4], traffic management [5], and sensor networks [6], as well as in emerging areas such as energy management [7, 8], autonomous vehicle navigation [9, 10, 11], and human–robot interaction [12].
Despite its wide adoption, implementing Nash strategies in real-world engineering systems faces significant challenges due to uncertainties, disturbances, delays, and actuator imperfections [13, 14, 15]. Extensions such as the $\varepsilon$-equilibrium relax optimality by allowing bounded deviations from Nash strategies [16], enabling numerical approximations [17, 18] and supporting further studies on sensitivity and robustness [13, 14, 19, 20, 21].
Robustness of Nash equilibrium to uncertainty has been studied from various perspectives. For example, model uncertainties in differential games were addressed in [13, 14, 21], where [13] designed a finite-horizon feedback Nash strategy for Linear Quadratic (LQ) games, [14] extended the scope to infinite-horizon settings, and [21] relaxed the LQ structure via weaker convexity assumptions. Other works addressed different uncertainty sources: [19] proposed adaptive control to mitigate measurement noise, and [20] analyzed the effect of cost parameter variations in discrete-time games.
@@
To address these effects, we propose a modified feedback policy that both mitigates performance loss and exploits available deviation information. While prior adaptive multi-agent methods exist [22, 23], they do not explicitly target compensation for execution-time control deviations of a specific player, which is addressed in this work.
@@ II Preliminaries
We focus on finite-horizon linear quadratic (LQ) games [24] with discrete-time dynamics
@@
We use Nash equilibrium to mean a strategy profile $U_{i}$, where no player can reduce its cost by a unilateral deviation, i.e. $J_{i}(x_{0},U_{i}^{\star},U_{-i}^{\star})\leq J_{i}(x_{0},U_{i},U_{-i}^{\star})$ [3].
@@
The FNE gains $\{K^{\star}_{i,k}\}$ are obtained by a backward Riccati recursion coupled across players [25]. Let $\tilde{P}_{i,N}=Q_{i,N}$ and define, for each $k\in\{0,\dots,N-1\}$ (computed backward),
@@
Under the standard assumptions $Q_{i},Q_{i,N}\succeq 0$, $R_{i}\succ 0$ and nonsingularity of the gain system in Definition 1 at each $k$, the backward sweep $k=N-1\to 0$ yields a unique FNE [24, Sec. 6].
@@
Under conditions 1-3, a unique finite-horizon Nash equilibrium exists. We assume these conditions henceforth [26].
@@ IV-A Compensated Feedback under Lagged Execution Error Dynamics
Inspired by the Disturbance-Based State Compensation (DBSC) approach [27], which is commonly used in traditional state-feedback control design, and given Assumption 1, we define the augmented state as $z_{k}:=[x_{k}^{\top}\ w_{k}^{\top}]^{\top}$.

### v3

@@ Abstract
Game-theoretic approaches and Nash equilibrium have been widely applied across various engineering domains. However, practical challenges such as disturbances, delays, and actuator limitations can hinder the precise execution of Nash equilibrium strategies. This work investigates the impact of such implementation imperfections on game trajectories and players’ costs in the context of a two-player finite-horizon linear quadratic (LQ) nonzero-sum game. Specifically, we analyze how small deviations by one player, measured or estimated at each stage, affect the state and cost function of the other player. To mitigate these effects, we propose an adjusted control policy that optimally compensates for the deviations under the stated information structure and can, under certain conditions, exploit them to improve performance. Rigorous mathematical analysis and proofs are provided, and the effectiveness of the proposed method is demonstrated through a representative numerical example.
@@ I Introduction
Game theory is a powerful analytical framework for modeling rational decision-making in complex multi-agent systems, particularly when interactions are strategic and agents’ decisions influence each other [shamma2020game, basar1999dynamic]. Such interactions often involve conflicting objectives, motivating the development of solution concepts to characterize equilibrium behavior. Among these, the Nash equilibrium is a well-established notion, defined as a set of strategies from which no participant can unilaterally deviate to achieve a better outcome [nash1950equilibrium].
Thanks to its solid theoretical foundation, Nash equilibrium has been widely applied in diverse engineering domains, including robotics [chiu2021encoding], traffic management [Oszczypała2023Nash], and sensor networks [Zheng2023game], as well as in emerging areas such as energy management [YANG2020216nashQ, abdolmohammadi2024sizing], autonomous vehicle navigation [nazari2024powertrain, bouzidi2025interaction, abtahi2025multi], and human–robot interaction [soltanian2025peerawarecostestimationnonlinear].
Despite its wide adoption, implementing Nash strategies in real-world engineering systems faces significant challenges due to uncertainties, disturbances, delays, and actuator imperfections [van-den=broek2023, AMATO2002507_guaranteeing, ma2024ϵ]. Extensions such as the $\varepsilon$-equilibrium relax optimality by allowing bounded deviations from Nash strategies [TANAKA1991413], enabling numerical approximations [NORTMANN20231760, alvarez2009urban] and supporting further studies on sensitivity and robustness [van-den=broek2023, AMATO2002507_guaranteeing, Jimenez01072006, tan2005sensitivity, hernandez2024ϵ].
Robustness of Nash equilibrium to uncertainty has been studied from various perspectives. For example, model uncertainties in differential games were addressed in [van-den=broek2023, AMATO2002507_guaranteeing, hernandez2024ϵ], where [van-den=broek2023] designed a finite-horizon feedback Nash strategy for Linear Quadratic (LQ) games, [AMATO2002507_guaranteeing] extended the scope to infinite-horizon settings, and [hernandez2024ϵ] relaxed the LQ structure via weaker convexity assumptions. Other works addressed different uncertainty sources: [Jimenez01072006] proposed adaptive control to mitigate measurement noise, and [tan2005sensitivity] analyzed the effect of cost parameter variations in discrete-time games.
@@
To address these effects, we propose a modified feedback policy that both mitigates performance loss and exploits available deviation information. While prior adaptive multi-agent methods exist [guerrero2021openloop, wang2022risk], they do not explicitly target compensation for execution-time control deviations of a specific player, which is addressed in this work.
@@ II Preliminaries
We focus on finite-horizon linear quadratic (LQ) games [Engwerda2009lq] with discrete-time dynamics
@@
We use Nash equilibrium to mean a strategy profile $U_{i}$, where no player can reduce its cost by a unilateral deviation, i.e. $J_{i}(x_{0},U_{i}^{\star},U_{-i}^{\star})\leq J_{i}(x_{0},U_{i},U_{-i}^{\star})$ [nash1950equilibrium].
@@
The FNE gains $\{K^{\star}_{i,k}\}$ are obtained by a backward Riccati recursion coupled across players [salizzoni2025bridging]. Let $\tilde{P}_{i,N}=Q_{i,N}$ and define, for each $k\in\{0,\dots,N-1\}$ (computed backward),
@@
Under the standard assumptions $Q_{i},Q_{i,N}\succeq 0$, $R_{i}\succ 0$ and nonsingularity of the gain system in Definition 1 at each $k$, the backward sweep $k=N-1\to 0$ yields a unique FNE [Engwerda2009lq, Sec. 6].
@@
Under conditions 1-3, a unique finite-horizon Nash equilibrium exists. We assume these conditions henceforth [LAA_Jungers_2013].
@@ IV-A Compensated Feedback under Lagged Execution Error Dynamics
Inspired by the Disturbance-Based State Compensation (DBSC) approach [rugh1996linear], which is commonly used in traditional state-feedback control design, and given Assumption 1, we define the augmented state as $z_{k}:=[x_{k}^{\top}\ w_{k}^{\top}]^{\top}$.